3400 dollars is placed in an account with an annual interest rate of 8.25%. How much will be in the account after 25 years, to the nearest cent?


1
Expert's answer
2022-03-17T10:10:09-0400

"A=S(1+r)^t" is the formula for an annual interest rate


A - the future amount

S - initial amount

r - rte of interest

t - time in years

Now plug in all variables into the equation

"A=3400(1+0.0825)^{25}\\=24670.419"

To the nearest cent would be rounding the answer to the nearest hundredth.

"24670.419=24670.42"